OpenStudy (amtran_bus):

How do you solve Log base sqrt 2 at 4?

OpenStudy (anonymous):

y = log (base a) x and a^y = x let a= b^2 so (b^2)^y =x if you take the sqrt, then b^y = sqrt(x) and log (base b) =sqrt(x)

OpenStudy (amtran_bus):